OpenStudy (anonymous):

find the equcation of plane through the points p(4,4,-4),q(2,-5,-5),r(1,2,2)

OpenStudy (loser66):

\[\vec {pq}=?\]

OpenStudy (loser66):

\[\vec{pr}=?\] then \[n=<\vec{pq}>.<\vec{pr}>\] \[\vec n\]and one point you can write out the equation of the required plane

OpenStudy (anonymous):

my answers was -56x+21y-23z=-48

OpenStudy (anonymous):

is wrong

OpenStudy (anonymous):

why

OpenStudy (anonymous):

\[\vec{PQ} = \vec{Q} - \vec{P}\]\[\vec{PR} = \vec{R} - \vec{P}\]\[\vec{n} = \vec{PQ} \times\vec{PR}\]

OpenStudy (anonymous):

i did that

OpenStudy (anonymous):

still wrong

OpenStudy (anonymous):

i find the n then i sub p0

OpenStudy (anonymous):

find the value

OpenStudy (anonymous):

pq=-2,-9,-1 pr=-3,-2,6

OpenStudy (anonymous):

pls check then i used matrix 3*3 matrix to find the n then i left with

OpenStudy (anonymous):

-56x+21y-23z

OpenStudy (anonymous):

then i sub p0 to get -48

OpenStudy (anonymous):

-56x + 15y - 23z + 72 = 0

OpenStudy (anonymous):

thanks